OpenStudy (anonymous):

[3f(x)+f(x)g(x)]=36 find f(2) suppose f and g are continuous functions such that g(2)=6

OpenStudy (anonymous):

\[3f(2)+f(2)*6=36\] \[f(2)[3+6]=36\] \[f(2)=\frac{ 36 }{ 9 }\]